# RealDex: Towards Human-like Grasping for Robotic Dexterous Hand

### v1

@@ Abstract
In this paper, we introduce RealDex, a pioneering dataset capturing authentic dexterous hand grasping motions infused with human behavioral patterns, enriched by multi-view and multimodal visual data. Utilizing a teleoperation system, we seamlessly synchronize human-robot hand poses in real time. This collection of human-like motions is crucial for training dexterous hands to mimic human movements more naturally and precisely. RealDex holds immense promise in advancing humanoid robot for automated perception, cognition, and manipulation in real-world scenarios. Moreover, we introduce a cutting-edge dexterous grasping motion generation framework, which aligns with human experience and enhances real-world applicability through effectively utilizing Multimodal Large Language Models. Extensive experiments have demonstrated the superior performance of our method on RealDex and other open datasets. The complete dataset and code will be made available upon the publication of this work.
@@
22footnotetext: These authors contribute equally to this work.11footnotetext: Corresponding author.
@@ 1 Introduction
Leveraging the RealDex dataset, we have developed a novel and effective method for dexterous grasping, which involves two key stages: grasp pose generation and motion synthesis, using only the point cloud data of objects captured by real vision sensor as input. Thanks to the dataset’s rich representation of human-like dexterous hand motions, our model can innately learn human behavior patterns, producing practical grasping motions. Recognizing the robust generalization capabilities of Multimodal Large Language Models (MLLMs) Gem (2023); GPT (2023), we have integrated an MLLM Selection Module. This module is adept at choosing the most natural, physically plausible, and human-like grasp pose from multiple generated options, enhancing the final motion synthesis. The inherent general knowledge within large models bolsters our method’s generalization capacity, ensuring effective performance even with unseen objects and thus significantly benefiting practical applications. We have conducted extensive experiments on RealDex and other open datasets, including dexterous grasping dataset and human grasping dataset, and also have tested our method on real robot hand. Quantitative and qualitative results demonstrate that our method outperforms others obviously for generation human-like practicable dexterous grasping motions.
@@ 3 RealDex Dataset
To advance research and practical applications of human-like robotic dexterous grasping in real-world environments, we introduce RealDex, the first extensive dataset that captures real dexterous hand grasping motions incorporating human behavioral patterns. RealDex features a diverse collection of 52 objects varying in scale, shape, and material, along with 2.6k sequences of grasping motions involving different initial object positions and grasping directions, and approximately 955k frames of raw visual data, including images and point clouds. We have divided our dataset into training, validation, and test sets, ensuring that each object appears exclusively in one of these subsets. The three sets contain 2114 grasping motion sequences for 40 objects, 245 grasping motions for 6 objects, and 271 grasping motions for 6 objects, respectively. Detailed information about the division will be provided in the supplementary materials.
We will delve into the specifics of RealDex, covering aspects such as the hardware setup of the capture system in Section 3.1, the calibration and synchronization of all sensors in Section 3.2, the data processing and annotation in Section 3.3, and the characteristics of RealDex in Section 3.4.
@@ 3.1 Hardware System Setup
Vision Capture System Much like human eyes, vision sensors are crucial for robots to perceive and understand their surroundings, thereby enabling appropriate action. Image-based and point cloud-based deep learning methods Qi et al. (2017a, b); Xu et al. (2023b); Zhu et al. (2021); Wang et al. (2021) have achieved significant breakthroughs in scene understanding, propelling the advancement of robotic perception systems. To support the practical deployment of dexterous hands, we have established a vision system equipped with four Azure Kinect RGB-D cameras with 15Hz capture frequency. These cameras are strategically positioned at four different locations, offering a quartet of perspectives focused on the operating desk, as illustrated in Figure 2. This setup provides RGB images and point clouds from multiple angles, not only aiding in delivering precise object pose annotations for our datasets, but also enabling research and practical applications of vision-based dexterous grasping under various visual configurations, including monocular-based, multi-view-based, point-based, multimodal-based approaches, etc.
Dexterous Manipulation System The dexterous manipulation system is designed to execute grasping operations akin to human actions. It features a UR10e robotic arm with 6 DoF and a tendon-driven Shadow Right Hand robot Sha (2005). This robot hand is equipped with 20 motors, enabling adduction and abduction movements across 24 degrees of freedom. Additionally, the dexterous hand is outfitted with over 100 sensors operating at up to 1 kHz.
Teleoperation System For the teleoperation system Tel (2022), users don a lightweight glove to intuitively control the Shadow Hand and robotic arm, allowing them to mimic natural human movements in real time. Human hand poses are seamlessly transmitted to the robot hand through the glove, vision tracking devices, and WiFi services. This system is utilized to gather data on human-like dexterous grasping motions, performed by various well-trained individuals.
@@ 3.2 Calibration and Synchronization
Camera-camera Calibration Camera intrinsics comes from manufacturer, while camera extrinsics are carefully measured through a two-stage process. Firstly, we use AprilTag Krogius et al. (2019) to roughly calculate extrinsics for cameras. Then, we apply the Iterative Closet Point method (ICP) Besl and McKay (1992); stytim (2023) between two neighbouring cameras to get more accurate alignment among point clouds from different views.
Camera-camera Synchronization The Kinect depth camera measures depth by timing the travel of emitted infrared light to objects and back. Interference from simultaneous infrared emissions by multiple cameras leads to depth inaccuracies, which we address by staggering camera sampling times and synchronizing auxiliary camera data streams with the main camera using aligned timestamps.
Robot-camera Calibration To synchronize the robotic hand’s motion with the vision system, we use hand-eye calibration IFL-CAMP (2021). This process starts with the Shadow Hand holding a calibration board, followed by moving it to over 30 locations for initial static transformation. We then refine this alignment by applying the Iterative Closest Point (ICP) Besl and McKay (1992) method between the robot’s mesh vertices and the main camera’s point cloud. This thorough calibration, as shown in Figure 3, ensures accurate alignment between the robotic and vision systems, enhancing data precision.
@@
Robot-camera Synchronization The robot’s timestamp, derived from the ROS TF tree, continuously updates the pose of each dexterous hand joint. To ensure precise time alignment, we match the initial camera frame’s point cloud with the robot’s mesh, reconstructed for each timestamp within a proximate time window. We then identify the robot frame most similar to the point cloud for synchronization. With this initial aligned timestamp established, we can extrapolate and align the remaining timestamps, ensuring a consistent and accurate correlation across the entire data set.
3.3 Data Processing
To enrich the diversity of data in RealDex, we have undertaken post-processing of the point clouds and object 6D poses, broadening the scope for research applications and facilitating more comprehensive studies.
Point Cloud Processing Beyond the raw RGB-D data captured by cameras, RealDex also offers 3D point clouds with more explicit geometric information, aiding point cloud-based dexterous grasping. To counteract the noise common in raw point clouds due to camera limitations, we apply filters like hidden point, statistical outlier, and radius outlier removal. Additionally, we merge point clouds from various perspectives into a singular comprehensive cloud using the cameras’ extrinsic parameters, supporting the creation of multi-view visual data algorithms.
Object 6D Pose Annotation To enhance the utility of our RealDex dataset, we have included precise mesh models and 6D poses for all featured objects. We use the commercial 3D scanner, the EinScan-Pro+ Ein (2018), to generate detailed 3D mesh models. We then refine these models by filling any holes to ensure completeness. Building upon the mesh models and the merged, comprehensive point cloud described earlier, we automatically annotate the 6D poses of objects for all grasping motion sequences using the ICP Besl and McKay (1992) method. To guarantee the high quality of these annotations, each object’s 6D pose has been manually inspected and adjusted against the point cloud. As Figure 3 shows, the accurate annotations can benefit more research of model-based grasping methods or object-tracking methods.
@@ 3.4 Characteristics
Human-like Grasping Motion In the data capture process of our dataset, the dexterous hand is operated via a teleoperation system directly controlled by a human hand. This unique approach ensures that all grasping motions in RealDex authentically mirror human behavioral patterns. In contrast, previous datasets only have static grasp poses derived from physical optimization, lacking dynamic motion sequences and any real human behavioral insights. The inclusion of realistic, human-like motion data in RealDex is profoundly significant for advancing intelligent dexterous hand development, imbued with human habits. This dataset holds the potential to propel the evolution of humanoid robotics.
Real Robotic Dexterous Poses Using a real dexterous robotic hand for data collection, our RealDex dataset accurately captures the ground truth of dexterous poses in every frame. This allows for the direct and effective application of models trained on RealDex in real-world scenarios. In contrast, previous datasets with synthesized grasp poses exhibit a substantial disparity from real-world data, leading to performance gap in translating models trained on such data into practical applications.
Rich Real Visual Data We have set up a vision capture system to obtain multi-view and multimodal visual data. This wealth of information is dexterous grasping for advancing research in vision-based dexterous grasping methods, which are typically more applicable and effective in real-world scenarios. The lack of genuine visual data in previous datasets significantly limits their practical utility.
@@ 4.1 Problem Formulation
Dexterous Hand Representation A dexterous hand is driven by revolute joints that only allow rotation around a fixed single axis. Hence, the pose of a dexterous hand is represented by the joint angle configuration, denoted as $\theta\in\mathbb{R}^{22}$. The global translation and orientation of the hand are represented as $\eta\in\mathbb{R}^{6}$, where the orientation is represented in angle-axis form. The pose of a robotic hand $\phi=(\theta,\eta)\in\mathbb{R}^{28}$ is the combination of joint angle configuration and global 6D pose of hand. The mesh representation of a dexterous hand can be generated by employing forward kinematics to calculate the global transformation for each segment. To represent the shape of hand mesh, a set of points is sampled, which we denote as $\mathbf{P}^{h}$.
Problem Formulation Given an observed object point cloud $\mathbf{P}^{o}$, we aim to generate the preferred grasping poses $\{\phi^{k}\}_{k=1}^{K}$ and synthesize the corresponding motion sequence $\Phi=\{\phi_{t}^{k}\}_{t=1}^{T}$ of dexterous hand to approach and grasp object for each grasping pose $\phi^{k}$.
@@ 4.2 Grasping Pose Generation
Pose Candidate Generation Inspired by the work in human grasping generation Hasson et al. (2019b); Jiang et al. (2021), we use a contact-aware generative model for dexterous grasp generation. This model uses a module to predict the contact map and a cVAE to learn a latent embedding grasping space. By sampling from the Gaussian variational distribution, the decoder of the cVAE is capable of generating a diverse array of candidate grasping poses $\{\phi^{k}\}_{k=1}^{N}$ for any input object point cloud. Subsequently, these poses undergo optimization to align with the predicted contact map, ensuring they are feasible. We then filter out unstable grasps that have insufficient contact with the object, and a human-like selection module is utilized to determine the optimal target poses $\{\phi^{k}\}_{k=1}^{K}$ for execution, where $K<N$.
Human-like Pose Selection Despite the acquisition of a diverse range of candidate grasping poses through consideration of the interaction between the object and the hand, this set of candidate poses fails to account for human grasping preferences and priors. Since MLLMs Gem (2023); GPT (2023) have already demonstrated their capability to encode rich world knowledge, we propose using MLLMs to discern the most natural, physically plausible, and human-like grasp poses from the candidate poses, thereby aligning it with human experience. Specifically, we begin by acquiring mesh representations of both the robotic hand and the object for each candidate grasping pose, followed by rendering the image depicting their interaction. Next, we prompt Gemini Gem (2023) using the specific prompt shown in Figure 4. The prompt emphasizes a comprehensive understanding of the grasping posture, highlighting the nuanced interactions between the fingers and the object. Gemini scores the rendered image, which contains intricate hand-object interaction details, in terms of naturalness, physical plausibility, human-likeness, and preference in order to select the target poses $\{\phi^{k}\}_{k=1}^{K}$.
@@ 5 Experiments
We developed and executed our algorithm using Python with PyTorch framework. Models were both trained and tested on an Ubuntu server, equipped with eight NVIDIA GeForce RTX 3090 GPU cards. We leveraged Gemini Gem (2023) for our MLLM selection module. More details for the training process, inference process, loss functions are introduced in the supplementary material. In the following, we first show the comparision results for vision-based grasping methods, and then evaluate each stage of our method in detail. Especially, test result on real robot is also demonstrated.
@@ 5.2 Evaluation for Grasping Pose Generation
Table 3: Comparison for grasp pose generation on three datasets.
@@
The self-intersection of robot hand and the intersection between hand and object are physically implausible, which are measured by the intersection volume (self.inter.vol) and the penetration distance (pene.dist). To measure the stability of grasping, we compute the simulation displacement (sim.disp.) following Tzionas et al. (2015); Hasson et al. (2019b), which is defined as the average displacement of the object when the hand is stationary and the object is subjected to gravity. These metrics provide a quantitative basis for determining the realism and viability of the grasp. We generate 10 grasp poses for each object in the test set for the evaluation. In addition, we also use the user score by rendering images for randomly selected 3 grasp poses of each object for more comprehensive evaluation.
@@ A.1 Point Cloud Denoising
Hardware limitations of RGBD cameras can introduce noise in the point clouds they generate, which may affect the accuracy of pose annotation. We refine our point cloud data by applying a statistical outlier removal filter Zhou et al. [2018]. This process involves analyzing each point’s average distance to its 20 nearest neighbors and excluding those points whose distance deviates by more than two standard deviations from the mean, effectively reducing noise.

### v2

@@ Abstract
In this paper, we introduce RealDex, a pioneering dataset capturing authentic dexterous hand grasping motions infused with human behavioral patterns, enriched by multi-view and multimodal visual data. Utilizing a teleoperation system, we seamlessly synchronize human-robot hand poses in real time. This collection of human-like motions is crucial for training dexterous hands to mimic human movements more naturally and precisely. RealDex holds immense promise in advancing humanoid robot for automated perception, cognition, and manipulation in real-world scenarios. Moreover, we introduce a cutting-edge dexterous grasping motion generation framework, which aligns with human experience and enhances real-world applicability through effectively utilizing Multimodal Large Language Models. Extensive experiments have demonstrated the superior performance of our method on RealDex and other open datasets. The dataset and associated code are available at https://4dvlab.github.io/RealDex_page/.
@@ 1 Introduction
Leveraging the RealDex dataset, we have developed a novel and effective method for dexterous grasping, which involves two key stages: grasp pose generation and motion synthesis, using only the point cloud data of objects captured by real vision sensor as input. Thanks to the dataset’s rich representation of human-like dexterous hand motions, our model can innately learn human behavior patterns, producing practical grasping motions. Recognizing the robust generalization capabilities of Multimodal Large Language Models (MLLMs) Google (2023); OpenAI (2023), we have integrated an MLLM Selection Module. This module is adept at choosing the most natural, physically plausible, and human-like grasp pose from multiple generated options, enhancing the final motion synthesis. The inherent general knowledge within large models bolsters our method’s generalization capacity, ensuring effective performance even with unseen objects and thus significantly benefiting practical applications. We have conducted extensive experiments on RealDex and other open datasets, including dexterous grasping dataset and human grasping dataset, and also have tested our method on real robot hand. Quantitative and qualitative results demonstrate that our method outperforms others obviously for generation human-like practicable dexterous grasping motions.
@@ 3 RealDex Dataset
To advance research and practical applications of human-like robotic dexterous grasping in real-world environments, we introduce RealDex, the first extensive dataset that captures real dexterous hand grasping motions incorporating human behavioral patterns. RealDex features a diverse collection of 52 objects varying in scale, shape, and material, along with 2.6k sequences of grasping motions involving different initial object positions and grasping directions, and approximately 955k frames of raw visual data, including images and point clouds. We use the commercial 3D scanner, the EinScan-Pro+ 3D (2018), to reconstruct detailed 3D mesh models for collected objects. We have divided our dataset into training, validation, and test sets, ensuring that each object appears exclusively in one of these subsets. The three sets contain 2114 grasping motion sequences for 40 objects, 245 grasping motions for 6 objects, and 271 grasping motions for 6 objects, respectively. Detailed information about the division will be provided in the supplementary materials.
We will delve into the specifics of RealDex, covering aspects such as the hardware setup of the capture system in Section 3.1, the calibration and synchronization of all sensors in Section 3.2, and the characteristics of RealDex in Section 3.3.
@@ 3.1 Hardware System Setup
Vision Capture System.
Much like human eyes, vision sensors are crucial for robots to perceive and understand their surroundings, thereby enabling appropriate action. Image-based and point cloud-based deep learning methods Qi et al. (2017a, b); Xu et al. (2023b); Zhu et al. (2021); Wang et al. (2021) have achieved significant breakthroughs in scene understanding, propelling the advancement of robotic perception systems. To support the practical deployment of dexterous hands, we have established a vision system equipped with four Azure Kinect RGB-D cameras with 15Hz capture frequency. These cameras are strategically positioned at four different locations, offering a quartet of perspectives focused on the operating desk, as illustrated in Figure 2. This setup provides RGB images and point clouds from multiple angles, not only aiding in delivering precise object pose annotations for our datasets, but also enabling research and practical applications of vision-based dexterous grasping under various visual configurations, including monocular-based, multi-view-based, point-based, multimodal-based approaches, etc.
Dexterous Manipulation System.
The dexterous manipulation system is designed to execute grasping operations akin to human actions. It features a UR10e robotic arm with 6 DoF and a tendon-driven Shadow Right Hand robot Sha (2005). This robot hand is equipped with 20 motors, enabling adduction and abduction movements across 24 degrees of freedom. Additionally, the dexterous hand is outfitted with over 100 sensors operating at up to 1 kHz.
Teleoperation System.
For the teleoperation system Tel (2022), users wear a lightweight glove to intuitively control the Shadow Hand and robotic arm, allowing them to mimic natural human movements in real time. Human hand poses are seamlessly transmitted to the robot hand through the glove, vision tracking devices, and WiFi services. This system is performed by well-trained individuals to collect data on human-like dexterous grasping motions.
@@ 3.2 Calibration and Synchronization
Camera-camera Calibration.
Camera intrinsics comes from manufacturer, while camera extrinsics are carefully measured through a two-stage process. Firstly, we use AprilTag Krogius et al. (2019) to roughly calculate extrinsics for cameras. Then, we apply the Iterative Closet Point method (ICP) Besl and McKay (1992); stytim (2023) between two neighbouring cameras to get more accurate alignment among point clouds from different views.
Camera-camera Synchronization.
The Kinect depth camera measures depth by timing the travel of emitted infrared light to objects and back. Interference from simultaneous infrared emissions by multiple cameras leads to depth inaccuracies, which we address by staggering camera sampling times and synchronizing auxiliary camera data streams with the main camera using aligned timestamps.
Robot-camera Calibration.
To synchronize the robotic hand’s motion with the vision system, we use hand-eye calibration IFL-CAMP (2021). This process starts with the Shadow Hand holding a calibration board, followed by moving it to over 30 locations for initial static transformation. We then refine this alignment by applying ICP method Besl and McKay (1992) between the robot’s mesh vertices and the main camera’s point cloud. This thorough calibration, as shown in Figure 3, ensures accurate alignment between the robotic and vision systems, enhancing data precision.
@@
Robot-camera Synchronization.
The robot’s timestamp, derived from the ROS TF tree, continuously updates the pose of each dexterous hand joint. To ensure precise time alignment, we match the initial camera frame’s point cloud with the robot’s mesh, reconstructed for each timestamp within a proximate time window. We then identify the robot frame most similar to the point cloud for synchronization. With this initial aligned timestamp established, we can extrapolate and align the remaining timestamps, ensuring a consistent and accurate correlation across the entire data set.
@@ 3.3 Characteristics
Human-like Grasping Motion.
In the data capture process of our dataset, the dexterous hand is operated via a teleoperation system directly controlled by a human hand. This unique approach ensures that all grasping motions in RealDex authentically mirror human behavioral patterns. In contrast, previous datasets only have static grasp poses derived from physical optimization, lacking dynamic motion sequences and any real human behavioral insights. The inclusion of realistic, human-like motion data in RealDex is profoundly significant for advancing intelligent dexterous hand development, imbued with human habits. This dataset holds the potential to propel the evolution of humanoid robotics.
Real Robotic Dexterous Poses.
Using a real dexterous robotic hand for data collection, our RealDex dataset accurately captures the ground truth of dexterous poses in every frame. This allows for the direct and effective application of models trained on RealDex in real-world scenarios. In contrast, previous datasets with synthesized grasp poses exhibit a substantial disparity from real-world data, leading to performance gap in translating models trained on such data into practical applications.
Rich Real Visual Data.
We have set up a vision capture system to obtain multi-view and multimodal visual data. This wealth of information is dexterous grasping for advancing research in vision-based dexterous grasping methods, which are typically more applicable and effective in real-world scenarios. The lack of genuine visual data in previous datasets significantly limits their practical utility.
@@ 4.1 Problem Formulation
Dexterous Hand Representation.
A dexterous hand is driven by revolute joints that only allow rotation around a fixed single axis. Hence, the pose of a dexterous hand is represented by the joint angle configuration, denoted as $\theta\in\mathbb{R}^{22}$. The global translation and orientation of the hand are represented as $\eta\in\mathbb{R}^{6}$, where the orientation is represented in angle-axis form. The pose of a robotic hand $\phi=(\theta,\eta)\in\mathbb{R}^{28}$ is the combination of joint angle configuration and global 6D pose of hand. The mesh representation of a dexterous hand can be generated by employing forward kinematics to calculate the global transformation for each segment. To represent the shape of hand mesh, a set of points is sampled, which we denote as $\mathbf{P}^{h}$.
Problem Formulation.
Given an observed object point cloud $\mathbf{P}^{o}$, we aim to generate the preferred grasping poses $\{\phi^{k}\}_{k=1}^{K}$ and synthesize the corresponding motion sequence $\Phi=\{\phi_{t}^{k}\}_{t=1}^{T}$ of dexterous hand to approach and grasp object for each grasping pose $\phi^{k}$.
@@ 4.2 Grasping Pose Generation
Pose Candidate Generation.
Inspired by the work in human grasping generation Hasson et al. (2019b); Jiang et al. (2021), we use a contact-aware generative model for dexterous grasp generation. This model uses a module to predict the contact map and a cVAE to learn a latent embedding grasping space. By sampling from the Gaussian variational distribution, the decoder of the cVAE is capable of generating a diverse array of candidate grasping poses $\{\phi^{k}\}_{k=1}^{N}$ for any input object point cloud. Subsequently, these poses undergo optimization to align with the predicted contact map, ensuring they are feasible. We then filter out unstable grasps that have insufficient contact with the object, and a human-like selection module is utilized to determine the optimal target poses $\{\phi^{k}\}_{k=1}^{K}$ for execution, where $K<N$.
Human-like Pose Selection.
Despite the acquisition of a diverse range of candidate grasping poses through consideration of the interaction between the object and the hand, this set of candidate poses fails to account for human grasping preferences and priors. Since MLLMs Google (2023); OpenAI (2023) have already demonstrated their capability to encode rich world knowledge, we propose using MLLMs to discern the most natural, physically plausible, and human-like grasp poses from the candidate poses, thereby aligning it with human experience. Specifically, we begin by acquiring mesh representations of both the robotic hand and the object for each candidate grasping pose, followed by rendering the image depicting their interaction. Next, we prompt Gemini Google (2023) using the specific prompt shown in Figure 4. The prompt emphasizes a comprehensive understanding of the grasping posture, highlighting the nuanced interactions between the fingers and the object. Gemini scores the rendered image, which contains intricate hand-object interaction details, in terms of naturalness, physical plausibility, human-likeness, and preference in order to select the target poses $\{\phi^{k}\}_{k=1}^{K}$.
@@ 5 Experiments
We developed and executed our algorithm using Python with PyTorch framework. Models were both trained and tested on an Ubuntu server, equipped with eight NVIDIA GeForce RTX 3090 GPU cards. We leveraged Gemini Google (2023) for our MLLM selection module. More details for the training process, inference process, loss functions are introduced in the supplementary material. In the following, we first show the comparision results for vision-based grasping methods, and then evaluate each stage of our method in detail. Especially, test result on real robot is also demonstrated.
@@ 5.2 Evaluation for Grasping Pose Generation
Table 3: Comparison for grasp pose generation on three datasets. D.G.N is the abbreviation of DexGraspNet.
The self-intersection of robot hand and the intersection between hand and object are physically implausible, which are measured by the intersection volume (s.i.vol.) and the penetration distance (p.dist.). To measure the stability of grasping, we compute the simulation displacement (sim.disp.) following Tzionas et al. (2015); Hasson et al. (2019b), which is defined as the average displacement of the object when the hand is stationary and the object is subjected to gravity. These metrics provide a quantitative basis for determining the realism and viability of the grasp. We generate 10 grasp poses for each object in the test set for the evaluation. In addition, we also use the user score by rendering images for randomly selected 3 grasp poses of each object for more comprehensive evaluation.
@@ Acknowledgements
Contribution Statement
* Yumeng Liu, Yaxun Yang and Youzhuo Wang contributed equally to this work.
$\dagger$ Yuexin Ma supervised the project.
@@ A.1 Point Cloud Denoising
To enhance the utility of our RealDex dataset, we have included precise mesh models and 6D poses for all featured objects. We automatically annotate the 6D poses of objects for all grasping motion sequences using the ICP method Besl and McKay [1992] and manually inspect and adjust them against the point cloud. Hardware limitations of RGBD cameras can introduce noise in the point clouds they generate, which may affect the accuracy of pose annotation. We apply a statistical outlier removal filter Zhou et al. [2018] to post-process raw point cloud data and merge point clouds from various viewpoints into a unified one. The denoising process involves analyzing each point’s average distance to its 20 nearest neighbors and excluding those points whose distance deviates by more than two standard deviations from the mean, effectively reducing noise.